Che Problem 7-15 Economic ordering quantity with safety stock [LO7-5] Diagnostic Supplies has expected sales of 84100 units per year, carrying costs of $5 per unit, and an ordering cost of $10 per order a. What is the economic ordering quantity? Ec b-1. What is the average inventory? b-2. What is the total carrying cost? < Prey 15 of 16 Next>
Answer #1

1

-/2(Annualusageinunits)(Ordercost) V Amualcarjn aos per uni) EOQ =

Square root of ((2*84100*10)/5)=580 units

2

EOQ/2 = 580/2 = 290units

3

total carrying cost

EOQ*carrying cost per unit =580*$5 per unit=$2900